# Towards Fair RAG: On the Impact of Fair Ranking in Retrieval-Augmented Generation

### v1

@@ 1 Introduction
In recent years, the concept of fair ranking has emerged as a critical concern in modern information access systems [11]. However, despite its significance, fair ranking has yet to be thoroughly examined in the context of retrieval-augmented generation (RAG) [28], a rapidly advancing trend in natural language processing (NLP) systems [26]. To understand why this is important, consider the RAG system in Figure 1, where a user asks a question about running shoes. A classic retrieval system might return several documents containing information from various running shoe companies. If the RAG system only selects the top two documents, then information from the remaining two relevant companies will not be relayed to the predictive model and will likely be omitted from its answer. The fair ranking literature refers to this situation as unfair because some relevant companies (i.e., in documents at position 3 and 4) receive less or no exposure compared to equally relevant company in the top position [11].
@@ 2 Background & Related Work
Retrieval-Augmented Generation. RAG, a specific type of retrieval-enhanced machine learning (REML) [53, 26], has been widely adopted in various domains, including language modeling [25], question-answering [21], personalization [45, 44, 27, 34], and recommendation [55]. Studies on the evaluation of RAG models have primarily focused on their effectiveness, including end-to-end performance [21, 17, 28] and the assessment of individual components [13, 41, 42]. However, little research has focused on evaluating fairness in RAG models, with the exception of recent work [48], which improved demographic diversity in human image generation by conditioning a generative model with externally retrieved images that help debias the generation process.
@@ 3 Experimental Methodology
To address the issue of unfairness in the rankings passed to the generator, we can convert a deterministic retriever into a stochastic retriever, which has been shown to provide fair rankings in expectation [9]. By sampling a ranking from a distribution of rankings predicted to be relevant, we smooth the expected exposure of different relevant items to be similar and, therefore, fairer (Appendix E). Because decisions are stochastic, the fairness and quality of stochastic retrieval is evaluated based on a sample of rankings. Similarly, since a sampled ranking is processed by a generator, we also compute the expected effectiveness by using sampled rankings. The complete evaluation pipeline of a RAG system with a stochastic retriever is illustrated in Figure 2.
@@ 3.1 Construction of a Test Collection with Utility Labels
We define an item’s worthiness by the additional utility (utility-gain) it provides to a language model (specifically, the generator in a RAG system) when used to solve a specific task as part of the augmentation process. To assess this utility-gain, each item in the corpus is individually supplied to the generator along with an input question. The utility-gain is then calculated as the difference between the utility of the augmented generator and that of a baseline language model without any information about the item. Formally, let $u_{i}$ denote the baseline string utility score from the vanilla language model prompted only with the input question, and let $u_{j}$ represent the utility score from the language model with a prompt augmented by the $j$’th item $d_{j}$ in the corpus. The item $d_{j}$ is considered useful if the utility-gain $\delta_{j}=u_{j}-u_{i}$ is positive, and not useful otherwise (see Appendix B).
Therefore, the item-level utility labels are designed to be both task- and generator-dependent, as the utility of each item varies depending on the task and the language model used. This labeling process also aligns with the principles of task-based information retrieval [24].
@@ 3.2 Fairness-Aware Stochastic Retriever
Given $n$ items in a corpus C, a vector of retrieval scores $\mathbf{s}\in\mathbb{R}^{n}$ can be obtained from $\mathcal{R}(q,\textrm{C})$, which can be used to generate a ranked list $L$. We then min-max normalize retrieval scores to be in [0, 1] in order to construct a multinomial distribution over items [4]. The probability of an item $d$ being selected as the $i$’th item in a new ranking $\sigma$ through Plackett-Luce sampling is given by
@@
where $L_{1:i-1}$ is the partial ranking up to position $i-1$, $\bar{\mathbf{s}}$ represents the normalized retrieval score vector, and $\bar{\mathbf{s}}_{d}$ is the normalized score of item $d$. Using this probability, we iteratively sample an item, set its probability to 0, renormalize the distribution, and repeat the process. The probability of generating a complete ranking is then given by the product of the placement probabilities for each item, i.e., $p(\sigma|q)=\prod^{n}_{i=1}p(\sigma_{i}|\sigma_{1:i-1})$.
This repeated sampling and renormalization process can be efficiently managed using the Gumbel-Softmax trick [16, 31], which enables the sampling of rankings to be performed at the speed of sorting [35]. To do so, for each sampling iteration, we draw $U_{i}\sim\text{Uniform}(0,1)$, followed by generating a Gumbel noise $G_{i}=-\log(-\log(U_{i}))$. The probability of each sampled ranking $\sigma_{i}$ is then obtained by sorting the items based on their perturbed scores $\tilde{\mathbf{s}}_{d_{i}}=\bar{\mathbf{s}}_{d_{i}}+G_{i}$.
@@ 3.2.1 Controlling the Level of Fairness
Adjusting the level of randomization directly controls the degree of item-fairness, aligning with our goal to observe how varying levels of fairness in rankings affect the ranking and generation quality of a RAG model. To obtain the controllability, we follow the work of Diaz et al. [9] and use a temperature parameter $\alpha$. We apply the scalar $\alpha$ to each value in the normalized score vector $\bar{\mathbf{s}}$ by raising each value to the power of $\alpha$ [^3] We normalized the values to the range of [1, 2] instead of [0, 1]. The addition of 1 effectively serves the same purpose as adjusting a real-numbered $\alpha$. We chose this range to allow for an integer-valued $\alpha$. This process is done before the scores are passed to the sampling policy. Therefore, the modified sampling distribution is thus defined as:
@@ 3.3 Evaluation
As mentioned in Section 3, because we are dealing with stochastic retrievers, we need to measure the expected behavior of the system. Let $\mathcal{S}(\mathbf{s},N,k)$ be the stochastic sampler that samples a set of $N$ rankings $\Sigma=[\sigma_{1},\sigma_{2},\cdots,\sigma_{N}]$, given the retrieval scores $\mathbf{s}$, where each ranking $\sigma_{i}$ is truncated to the size of $k$. From each ranking, we can get an output $\hat{y}_{i}=\mathcal{G}(\phi_{p}(x,\sigma_{i}))$. With an arbitrary fairness metric $\mu_{f}(\Sigma)$ and a ranking quality metric $\mu_{r}(\Sigma)$ that takes a set of rankings as an input, we can measure the degree of fairness and ranking quality of the sampled rankings. Similarly, an arbitrary string utility metric $\mu_{u}(y,\hat{y}_{i})$, such as ROUGE, can be used to assess an expected effectiveness of a RAG system by calculating the average of the $N$ metric scores.
@@ 3.3.1 Expected Exposure in the Context of Machine Users
Given this machine user browsing model and a mapping from item index to its rank denoted as $\bar{\sigma}_{d}$, a system exposure for each item $d$ is calculated as
$\mathbf{\epsilon}_{d}=\sum_{\sigma\in S_{n}}\pi(\sigma|q)\text{MU}(\bar{\sigma}_{d})$ (5)
@@ 3.3.2 Expected Utility
Given the set of $N$ sampled rankings $\Sigma$, we individually augment the generator with each ranking $\sigma_{i}$, resulting in $N$ outputs from the generator. The utility of these outputs is then measured using an arbitrary string utility metric $\mu_{u}$. To determine the anticipated utility of a RAG model with fair rankings—represented by the tuple of a stochastic ranking sampler $\mathcal{S}$ and a generator $\mathcal{G}$—we calculate the expected utility (EU) of the RAG system given an instance $x$.
$\text{EU}(\langle\mathcal{S},\mathcal{G}\rangle|x)=\mathbb{E}_{\sigma\sim\mathcal{S}}[\mu_{u}(y,\hat{y}_{\sigma})]=\sum_{\sigma\in S_{n}}p(\sigma|q)\mu_{u}(y,\hat{y}_{\sigma})\approx\frac{1}{N}\sum^{N}_{i=1}\mu_{u}(y,\hat{y}_{i})$ (7)
where $\hat{y}_{\sigma}$ is the prediction of a system given the ranking $\sigma$, $S_{n}$ is the symmetric group of a ranked list $L$ from the deterministic retriever $\mathcal{R}$, and $\sum_{\sigma\in S_{n}}p(\sigma|q)=1$.
@@ 4 Experiment setup
Figure 3: Effect of a temperature parameter $\alpha$ on the disparity of rankings, in the LaMP task 4, a text generation task. The RAG model is configured with the Contriever and Flan-T5-Base. Each data point represents the normalized EE-D of each run of the experiment (i.e., one query –>$N$ sampled rankings –> $\overline{\text{EE-D}}$ of the $N$ rankings).
We used BM25 (lexical retriever) [39], SPLADE (learned sparse retriever), and Contriever (bi-encoder dense retriever) [20] as deterministic retrievers providing retrieval scores to base the sampling on. These models represent commonly used retrievers in the RAG literature [26]. Sampling size was set to $N=100$, and truncation size was set to $k=5$.
For generation models, we use Flan-T5-Small, Flan-T5-Base, and Flan-T5-XXL [7]. For decoding strategy, beam size is set to 4, and no sampling strategy is used. This is to ensure that stochasticity is only introduced to the retriever for controlled experiments. Full implementation details can be found in Appendix K. With the three base retrievers and three generators, we configure nine different RAG models and evaluate them on the seven LaMP tasks.
We repeat the experiments with four different temperature parameters $\alpha=1,2,4,8$, which allows us to assess the utility of the RAG models with different levels of item-fairness. From Figure 3, we observe how effectively $\alpha$, described in the Equation 2, controls the disparity of rankings. For example, when $\alpha$ is set to 4, we usually obtain a set of sampled rankings with $\overline{\text{EE-D}}$ mostly in the range of [0.5, 0.8], and when $\alpha$ is set to 8, we often get a set of sampled rankings with $\overline{\text{EE-D}}=1$. Refer to Appendix D to see the full description of the effect of $\alpha$ on the other metrics.
@@ 5 Results
The leftmost side of the lines, where $\overline{\text{EE-D}}=0$, represents the performance of a uniformly random ranking policy. At this point, the measured ranking quality should approximate the proportion of positively labeled items in the corpus, which is 31% based on data statistics. This is notably higher than in non-RAG (human-user) settings, where the percentage of relevant documents is typically much smaller, resulting in a $\overline{\text{EE-R}}$ value near 0 [9].
@@ 6 Discussion and Conclusion
Conclusion. This study highlights the impact of fair rankings on both the ranking and generation quality of RAG systems. Through the extensive analysis, we show that fairer RAG systems not only maintain high generation quality but can also outperform traditional RAG models, challenging the notion of a strict tradeoff between fairness and effectiveness. Our findings provide valuable insights for developing responsible and equitable RAG systems and pave the way for future research in fair ranking and retrieval-augmented generation. We encourage future researchers to extend this work by incorporating graded or missing judgments and exploring the different notions of fairness in RAG systems, and ultimately advancing the field of trustworthy RAG systems research.
@@ Appendix A Notation
$\sigma_{i}$ sampled ranking $\in\Sigma$
$\phi_{p}(x,\sigma_{i})$ prompt generation function
$\overline{x}_{i}$ prompt returned by $\phi_{p}(x,\sigma_{i})$
$\mathcal{G}(\overline{x}_{i})$ language model
$\hat{y}_{i}$ predicted output from $\mathcal{G}(\overline{x}_{i})$
@@
$\mu_{u}(y,\hat{y}_{i})$ string utility metric
@@ Appendix B Labeling Procedure
1:$\mathcal{D}=\{(x_{1},y_{1}),(x_{2},y_{2}),\cdots,(x_{d},y_{d})\}$
@@
7:         $\delta_{j}\leftarrow(u_{j}-u_{i})$triangleright extra utility gained from the augmentation
@@ C.1 Normalization of EE-D
For upper bound, recall that the $\mathbf{\epsilon}$ is computed based on samples of rankings from a stochastic policy $\pi$. For relevance-independent browsing models (e.g., UM), all rankings $\sigma\in S_{n}$ have identical exposure norms ${\left\lVert\mathbf{\epsilon}^{\sigma}\right\rVert}_{2}^{2}$, where $\mathbf{\epsilon}^{\sigma}$ is the exposure of items for ranking sampled from $\pi$. Then,
${\left\lVert\mathbf{\epsilon}\right\rVert}_{2}^{2}$ $={\left\lVert\mathbb{E}_{\sigma\sim\pi}[\mathbf{\epsilon}^{\sigma}]\right\rVert}_{2}^{2}$ (9)
$\leq\mathbb{E}_{\sigma\sim\pi}[{\left\lVert\mathbf{\epsilon}^{\sigma}\right\rVert}_{2}^{2}]$ (10)
$=\mathbb{E}_{\sigma\sim\pi}[{\left\lVert\bar{\mathbf{\epsilon}}\right\rVert}_{2}^{2}]={\left\lVert\bar{\mathbf{\epsilon}}\right\rVert}_{2}^{2}$ (11)
@@
With UM, the upper bound of EE-D, ${\left\lVert\bar{\mathbf{\epsilon}}\right\rVert}_{2}^{2}$ becomes $\sum_{i=1}^{n}{\text{UM}(i)}^{2}=k$. Therefore, per query $q$, we calculate a normalized EE-D

### v2

@@ 1 Introduction
In recent years, the concept of fair ranking has emerged as a critical concern in modern information access systems [12]. However, despite its significance, fair ranking has yet to be thoroughly examined in the context of retrieval-augmented generation (RAG) [29, 1], a rapidly advancing trend in natural language processing (NLP) systems [27]. To understand why this is important, consider the RAG system in Figure 1, where a user asks a question about running shoes. A classic retrieval system might return several documents containing information from various running shoe companies. If the RAG system only selects the top two documents, then information from the remaining two relevant companies will not be relayed to the predictive model and will likely be omitted from its answer. The fair ranking literature refers to this situation as unfair because some relevant companies (i.e., in documents at position 3 and 4) receive less or no exposure compared to equally relevant company in the top position [12].
@@ 2 Background & Related Work
Retrieval-Augmented Generation. RAG, a specific type of retrieval-enhanced machine learning (REML) [54, 27], has been widely adopted in various domains, including language modeling [26], question-answering [22], personalization [46, 45, 28, 35], and recommendation [56]. Studies on the evaluation of RAG models have primarily focused on their effectiveness, including end-to-end performance [22, 18, 29] and the assessment of individual components [14, 42, 43]. However, little research has focused on evaluating fairness in retrieval-enhanced generation models, with the exception of recent work [49], which improved demographic diversity in human image generation by conditioning a generative model with externally retrieved images that help debias the generation process.
@@ 3 Experimental Methodology
To address the issue of unfairness in the rankings passed to the generator, we can convert a deterministic retriever into a stochastic retriever, which can, in expectation, provide fair rankings [10]. By sampling a ranking based on its quality to users—in this case generators—the expected exposure of different relevant items becomes similar and, therefore, fairer (Appendix E). Because decisions are stochastic, the fairness and quality of stochastic retrieval is evaluated based on a sample of rankings. Similarly, since a sampled ranking is processed by a generator, we also compute the expected generator effectiveness over sampled rankings. The complete evaluation pipeline of a RAG system with a stochastic retriever is illustrated in Figure 2.
@@ 3.1 Construction of a Test Collection with Utility Labels
We define an item’s worthiness by the marginal gain in utility (utility-gain) it provides to a language model (specifically, the generator in a RAG system) when used to solve a specific task as part of the augmentation process. To assess this utility-gain, each item in the corpus is individually supplied to the generator along with an input question. The utility-gain is then calculated as the difference between the utility of the augmented generator and that of a baseline language model without any information about the item. Formally, let $u_{i}$ denote the baseline string utility score from the vanilla language model prompted only with the input question, and let $u_{j}$ represent the utility score from the language model with a prompt augmented by the $j$’th item $d_{j}$ in the corpus. The item $d_{j}$ is considered useful if the utility-gain $\delta_{j}=u_{j}-u_{i}$ is positive, and not useful otherwise (see Appendix B).
Therefore, the item-level utility labels are designed to be both task- and generator-dependent, as the utility of each item varies depending on the task and the language model used. This labeling process also aligns with the principles of task-based information retrieval, where, in the context of human searchers, document utility may vary on how the user expects to use the document [25].
@@ 3.2 Fairness-Aware Stochastic Retriever
Given $n$ items in a corpus C, a vector of retrieval scores $\mathbf{s}\in\mathbb{R}^{n}$ can be obtained from $\mathcal{R}(q,\textrm{C})$, which can be used to generate a ranked list $L$. We then min-max normalize retrieval scores to be in [0, 1] in order to construct a multinomial distribution over items [5]. The probability of an item $d$ being selected as the $i$’th item in a new ranking $\pi$ through Plackett-Luce sampling is given by
@@
where $L_{1:i-1}$ is the partial ranking up to position $i-1$, $\bar{\mathbf{s}}$ represents the normalized retrieval score vector, and $\bar{\mathbf{s}}_{d}$ is the normalized score of item $d$. Using this probability, we iteratively sample an item, set its probability to 0, renormalize the distribution, and repeat the process. The probability of generating a complete ranking is then given by the product of the placement probabilities for each item, i.e., $p(\pi|q)=\prod^{n}_{i=1}p(\pi_{i}|\pi_{1:i-1})$.
This repeated sampling and renormalization process can be efficiently managed using the Gumbel-Softmax trick [17, 32], which enables the sampling of rankings to be performed at the speed of sorting [36]. To do so, for each sampling iteration, we draw $U_{i}\sim\text{Uniform}(0,1)$, followed by generating a Gumbel noise $G_{i}=-\log(-\log(U_{i}))$. The probability of each sampled ranking is then obtained by sorting the items based on their perturbed scores $\tilde{\mathbf{s}}_{d_{i}}=\bar{\mathbf{s}}_{d_{i}}+G_{i}$.
@@ 3.2.1 Controlling the Level of Fairness
Adjusting the level of randomization directly controls the degree of item-fairness, aligning with our goal to observe how varying levels of fairness in rankings affect the ranking and generation quality of a RAG model. To obtain the controllability, we follow the work of Diaz et al. [10] and use a fairness control parameter $\alpha$. We apply the scalar $\alpha$ to each value in the normalized score vector $\bar{\mathbf{s}}$ by raising each value to the power of $\alpha$ [^3] We normalized the values to the range of [1, 2] instead of [0, 1]. The addition of 1 effectively serves the same purpose as adjusting a real-numbered $\alpha$. We chose this range to allow for an integer-valued $\alpha$. This process is done before the scores are passed to the sampling policy. Therefore, the modified sampling distribution is thus defined as:
@@ 3.3 Evaluation
As mentioned in Section 3, because we are dealing with stochastic retrievers, we need to measure the expected behavior of the system. Let $\mathcal{S}(\mathbf{s},N,k)$ be the stochastic sampler that samples a set of $N$ rankings $\sigma=\{\pi\}$, given the retrieval scores $\mathbf{s}$, where each ranking $\pi$ is truncated to the size of $k$. From each ranking, we can get an output $\hat{y}=\mathcal{G}(\phi_{p}(x,\pi))$. With an arbitrary fairness metric $\mu_{f}(\sigma)$ and a ranking quality metric $\mu_{r}(\sigma)$ that takes a set of rankings as an input, we can measure the degree of fairness and ranking quality of the sampled rankings. Similarly, an arbitrary string utility metric $\mu_{u}(y,\hat{y})$, such as ROUGE, can be used to assess an expected effectiveness of a RAG system by calculating the average of the $N$ metric scores.
@@ 3.3.1 Expected Exposure in the Context of Machine Users
Given this machine user browsing model and a mapping from item index to its rank denoted as $\bar{\pi}_{d}$, a system exposure for each item $d$ is calculated as
$\mathbf{\epsilon}_{d}=\sum_{\pi\in S_{n}}p(\pi|q)\text{MU}(\bar{\pi}_{d})$ (5)
@@ 3.3.2 Expected Utility
Given the set of $N$ sampled rankings $\sigma$, we individually augment the generator with each ranking $\pi\in\sigma$, resulting in $N$ outputs from the generator. The utility of these outputs is then measured using an arbitrary string utility metric $\mu_{u}$. To determine the anticipated utility of a RAG model with fair rankings—represented by the tuple of a stochastic ranking sampler $\mathcal{S}$ and a generator $\mathcal{G}$—we calculate the expected utility (EU) of the RAG system given an instance $x$.
$\text{EU}(\langle\mathcal{S},\mathcal{G}\rangle|x)=\mathbb{E}_{\pi\sim\mathcal{S}}[\mu_{u}(y,\hat{y}_{\pi})]=\sum_{\pi\in S_{n}}p(\pi|q)\mu_{u}(y,\hat{y}_{\pi})\approx\frac{1}{N}\sum_{\pi\in\sigma}\mu_{u}(y,\hat{y}_{\pi})$ (7)
where $\hat{y}_{\pi}$ is the prediction of a system given the ranking $\pi$, $S_{n}$ is the symmetric group of a ranked list $L$ from the deterministic retriever $\mathcal{R}$, and $\sum_{\pi\in S_{n}}p(\pi|q)=1$.
@@ 4 Experiment setup
Figure 3: Effect of a fairness control parameter $\alpha$ on the disparity of rankings, in the LaMP task 4, a text generation task. The RAG model is configured with the Contriever and Flan-T5-Base. Each data point represents the normalized EE-D of each run of the experiment (i.e., one query –>$N$ sampled rankings –> $\overline{\text{EE-D}}$ of the $N$ rankings).
We use BM25 (lexical retriever) [40], SPLADE (learned sparse retriever), and Contriever (bi-encoder dense retriever) [21] as deterministic retrievers providing retrieval scores to base the sampling on. These models represent commonly used retrievers in the RAG literature [27]. We use a sampling size of $N=100$ and a truncation size of $k=5$.
For generation models, we use Flan-T5-Small, Flan-T5-Base, and Flan-T5-XXL [8]. For decoding strategy, beam size is set to 4, and no sampling strategy is used. This is to ensure that stochasticity is only introduced to the retriever for controlled experiments. Full implementation details can be found in Appendix K. With the three base retrievers and three generators, we configure nine different RAG models and evaluate them on the seven LaMP tasks. Utility measurement of the generated strings followed the metrics used in the LaMP paper.
We repeat the experiments with four different fairness control parameters $\alpha=1,2,4,8$, which allows us to assess the utility of the RAG models with different levels of item-fairness. From Figure 3, we observe how effectively $\alpha$, described in the Equation 2, controls the disparity of rankings. For example, when $\alpha$ is set to 4, we usually obtain a set of sampled rankings with $\overline{\text{EE-D}}$ mostly in the range of [0.5, 0.8], and when $\alpha$ is set to 8, we often get a set of sampled rankings with $\overline{\text{EE-D}}=1$. Refer to Appendix D to see the full description of the effect of $\alpha$ on the other metrics.
@@ 5 Results
The leftmost side of the lines, where $\overline{\text{EE-D}}=0$, represents the performance of a uniformly random ranking policy. At this point, the measured ranking quality should approximate the proportion of positively labeled items in the corpus, which is 31% based on data statistics (Appendix J). This is notably higher than in non-RAG (human-user) settings, where the percentage of relevant documents is typically much smaller, resulting in a $\overline{\text{EE-R}}$ value near 0 [10].
@@ 6 Discussion and Conclusion
Conclusion. This study highlights the impact of fair rankings on both the ranking and generation quality of RAG systems. Through the extensive analysis, we show that fairer RAG systems not only maintain high generation quality but can also outperform traditional RAG models, challenging the notion of a strict tradeoff between fairness and effectiveness. Our findings provide valuable insights for developing responsible and equitable RAG systems and pave the way for future research in fair ranking and retrieval-augmented generation. In future work, we hope to extend this framework to consider graded or missing judgments and exploring the different notions of fairness in RAG systems, ultimately advancing the field of trustworthy RAG systems research.
@@ Appendix A Notation
$\pi$ sampled ranking $\in\sigma$
$\phi_{p}(x,\pi)$ prompt generation function
$\overline{x}$ prompt returned by $\phi_{p}(x,\pi)$
$\mathcal{G}(\overline{x})$ language model
$\hat{y}$ predicted output from $\mathcal{G}(\overline{x})$
@@
$\mu_{u}(y,\hat{y})$ string utility metric
@@ Appendix B Labeling Procedure
1:$\mathcal{D}=\{(x_{1},y_{1}),(x_{2},y_{2}),\cdots,(x_{T},y_{T})\}$triangleright dataset of size $T$
@@
7:         $\delta_{j}\leftarrow(u_{j}-u_{i})$triangleright marginal utility gained from the augmentation
@@ C.1 Normalization of EE-D
For upper bound, recall that the $\mathbf{\epsilon}$ is computed based on samples of rankings from a stochastic policy. For relevance-independent browsing models (e.g., MU), all rankings $\pi\in S_{n}$ have identical exposure norms ${\left\lVert\mathbf{\epsilon}^{\pi}\right\rVert}_{2}^{2}$, where $\mathbf{\epsilon}^{\pi}$ is the exposure of items for ranking $\pi$ sampled from the stochastic policy. Then,
${\left\lVert\mathbf{\epsilon}\right\rVert}_{2}^{2}$ $={\left\lVert\mathbb{E}_{\pi}[\mathbf{\epsilon}^{\pi}]\right\rVert}_{2}^{2}$ (9)
$\leq\mathbb{E}_{\pi}[{\left\lVert\mathbf{\epsilon}^{\pi}\right\rVert}_{2}^{2}]$ (10)
$=\mathbb{E}_{\pi}[{\left\lVert\bar{\mathbf{\epsilon}}\right\rVert}_{2}^{2}]={\left\lVert\bar{\mathbf{\epsilon}}\right\rVert}_{2}^{2}$ (11)
@@
With MU, the upper bound of EE-D, ${\left\lVert\bar{\mathbf{\epsilon}}\right\rVert}_{2}^{2}$ becomes $\sum_{i=1}^{n}{\text{MU}(i)}^{2}=k$. Therefore, per query $q$, we calculate a normalized EE-D